# Chongkai Gao1, Zhehao Cai1, Jiayu Luo1, Lin Shao1

### v2

@@ I Introduction
Dexterous grasping is crucial in robotics as the first step in executing complex manipulation tasks. However, quickly obtaining a high-quality and diverse set of grasps remains challenging for dexterous robotic hands due to their high degrees of freedom and the complexities involved in achieving stable, precise grasps. Researchers have developed several optimization-based methods to address this challenge  [1, 2, 3, 4, 5]. Some of these methods, however, often focus on fingertip point contact, relying on complete object shape, and require significant computational time to optimize. As a result, data-driven grasp generation methods have gained attention. These methods aim to solve the grasping problem using learning-based techniques. We can broadly categorize them into two types: those that utilize robot-centric representations, such as wrist poses and joint values [6, 7, 8], and those that rely on object-centric representations, such as contact points [9, 10, 11] or contact maps [12, 13, 14, 15].
Robot-centric representations (e.g., joint values), as used in methods like UniDexGrasp++ [8], directly map observation to control commands, enabling fast inference. However, these methods often exhibit low sample efficiency. Furthermore, these approaches struggle to generalize across different robotic embodiments, as the learned mappings are specific to the training data and do not quickly adapt to new robot designs or geometries. Object-centric representations (e.g., key points, contact points, affordances) effectively capture the geometry and contacts of objects, allowing for generalization across different shapes and robots, as demonstrated by methods like UniGrasp [9] and GenDexGrasp [12]. However, these methods are often less efficient as they typically require an additional optimization step—such as solving fingertip inverse kinematics (IK) or fitting the predicted contact maps under penetration-free and joint limit constraints to translate the object-centric representation into actionable robot commands. This optimization process is time-consuming due to its complexity and nonconvexity [16, 17, 18].
@@
Full-hand
Contact
@@
Given the point clouds of both an open robotic hand and the object, our network architecture predicts the $\mathcal{D(R,O)}$ representation. This matrix encodes the relative distances between the point clouds of the object and the robotic hand in the desired grasping pose. Using this representation, we apply a multilateration method [19] to estimate the robot’s point cloud at the predicted pose, allowing us to compute the 6D pose of each hand link in the world frame and ultimately determine the joint configurations. To encode robotic hands, we propose a configuration-invariant pretraining method that learns the inherent alignment between various hand configurations, promoting grasp generation performance and cross-embodiment generalization. We validate the effectiveness of our approach through extensive experiments in both simulation and real-world settings. Our model achieves an average success rate of 87.53% in simulation across three dexterous robotic hands and in real-robot experiments, demonstrating its robustness and versatility.
@@ II-A Learning-based Robotic Dexterous Grasping
Another line of research focuses on robot-centric approaches, directly inferring robot poses [20] or joint angles [8]. These approaches are generally easier to execute as they directly infer joint values for dexterous hands. Reinforcement learning algorithms commonly generate configurations in high-dimensional action spaces and often train in simulation, which can be sample-inefficient. The sim-to-real gap [21, 22] further complicates policy transfer, and some methods only report results in simulation without real-world validation. Some works propose leveraging object-centric features [23], human grasp priors [24], or hand-object interaction features [25] to accelerate the policy generation process and address these issues. Robot-centric policy generation approaches often struggle to generalize across different robotic embodiments. To combine the strengths of both approaches, we propose learning the interaction relationship between the robot and object as a representation for dexterous robotic grasping, achieving robust real-world grasping performance and cross-embodiment generalization.
@@ III Method
III-A Configuration-Invariant Pretraining with Contrastive Learning
Learning dexterous grasping involves understanding the spatial relationships between the robot hand and the object. The objective is to match the robot hand in a specific configuration with the object. However, this matching process is challenging because the local geometric features of a point in the open-hand configuration may not align with those in the grasp configuration due to huge variations during articulation.
@@ III-A Configuration-Invariant Pretraining with Contrastive Learning
Specifically, for each robot hand, we begin by uniformly sampling points on the surface of each link at the canonical pose, storing the resulting point clouds denoted as $\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}$, where $N_{\ell}$ is the number of links. We define a point cloud forward kinematics model, $\text{FK}\left(q,\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}\right)$ to map joint configurations to point clouds at new poses. For example, given a close-hand $q_{\mathcal{A}}$ and an open-hand configuration $q_{\mathcal{B}}$, where the wrist pose is the same or nearly identical, we obtain two point clouds $\mathbf{P}^{\mathcal{A}},\mathbf{P}^{\mathcal{B}}\in\mathbb{R}^{N_{\mathcal{R}}\times 3}$, representing these two joint configurations. Here, $N_{\mathcal{R}}$ is the number of points in the robot point cloud, set to 512 in practice.
@@ III-B $\mathcal{D(R,O)}$ Prediction
Denote an open-hand configuration as $q_{\text{init}}$, of which the wrist pose can be either user-specified or randomly generated. Let the robot point cloud under $q_{\text{init}}$ be $\mathbf{P}^{\mathcal{R}}=\text{FK}\left(q_{\text{init}},\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}\right)\in\mathbb{R}^{N_{\mathcal{R}}\times 3}$, and the object point cloud be $\mathbf{P}^{\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times 3}$, where $N_{\mathcal{O}}$ represents the number of points in the object point cloud, also set to 512 in practice. The objective of our neural network is to predict the point-to-point distance matrix $\mathcal{D(R,O)}\in\mathbb{R}^{N_{\mathcal{R}}\times N_{\mathcal{O}}}$.
Point Cloud Feature Extraction We begin by extracting point cloud embeddings using two encoders, $f_{\theta_{\mathcal{R}}}(\mathbf{P}^{\mathcal{R}})$ and $f_{\theta_{\mathcal{O}}}(\mathbf{P}^{\mathcal{O}})$, which share the same architecture. Specifically, we use a modified DGCNN [27] to better capture local structures and integrate global information (See Appendix -F1). The robot encoder is initialized with pretrained parameters, using the method described in Sec. III-A, and remains frozen during training. These encoders extract point-wise features, $\bm{\phi}^{\mathcal{R}}$ and $\bm{\phi}^{\mathcal{O}}$ from the robot and object point clouds:
@@
To establish correspondences between the robot and object features, we apply two multi-head cross-attention transformers [28] (See Appendix -F2), $g_{\theta_{\mathcal{R}}}(\bm{\phi}^{\mathcal{R}},\bm{\phi}^{\mathcal{O}})$ and $g_{\theta_{\mathcal{O}}}(\bm{\phi}^{\mathcal{O}},\bm{\phi}^{\mathcal{R}})$. These transformers integrate the relationships between the two feature sets, embedding correspondence information. This process maps the robot and object features to two sets of correlated features, $\bm{\psi}^{\mathcal{R}}$ and $\bm{\psi}^{\mathcal{O}}$:
@@
where $\sigma$ denotes the softplus function, and $\mathcal{N}_{\theta}$ is an MLP, which takes in the feature of $\mathbb{R}^{N_{\mathcal{O}}\times(2D+2d)}$ and outputs a positive number (See Appendix -F3). By calculating on all $(\widehat{\bm{\psi}}_{i}^{\mathcal{R}},\widehat{\bm{\psi}}_{j}^{\mathcal{O}})$ pairs, we obtain the complete $\mathcal{D(R,O)}$ representation:
@@ III-C Grasp Configuration Generation from $\mathcal{D(R,O)}$
As shown in [31], this problem has a closed-form solution, and by using the implementation from [30], we can directly compute $p_{i}^{\prime\mathcal{R}}$. Repeating this process for each row of $\mathcal{D(R,O)}$ yields the complete predicted robot point cloud $\mathbf{P}^{\mathcal{P}}$ in the grasp pose. In 3D space, we can determine a point’s position by measuring its relative distances to just three other points. Our $\mathcal{D(R,O)}$ representation provides $N_{\mathcal{O}}(=512)$ relative distances, enhancing robustness to prediction errors.
@@
where $\mathbf{x}_{i}$ and $\mathbf{R}_{i}$ represent the translation and rotation of the $i$-th link, respectively.
Joint Configuration Optimization After predicting the 6D pose for each link, our objective is to optimize the joint values to align the translation of each link with the predicted result. Starting from an initial value $q_{init}$, we iteratively solve the following optimization problem using CVXPY [32]:
@@
In each iteration, the system computes the delta joint values $\delta\bm{q}$ by minimizing the objective function and updates the joint values as $\bm{q}\leftarrow\bm{q}+\delta\bm{q}$. Here, $\mathbf{x}_{i}$ represents the current link translation, $[\bm{q}_{min},\bm{q}_{max}]$ denotes the joint limits, and $\varepsilon_{q}=0.5$ is the maximum allowable step size. The optimization process can be efficiently parallelized, typically achieving convergence within one second, even for a 6+22 DoF ShadowHand.
@@ III-D Loss Function
where $\lambda_{\mathcal{D}}$, $\lambda_{\mathcal{T}}$, $\lambda_{\mathcal{P}}$, $\lambda_{KL}$ are hyperparameters for loss weights.The superscript ‘GT’ refers to the ground truth annotations. $\mathcal{N}(0,I)$ is a standard Guassian distribution, and $\mathbf{P}^{\mathcal{T}}$ is the robot point cloud under the $\bm{\mathcal{T}^{*}}$ described in  III-C. $\mathcal{L_{\text{P}}}$ computes the sum of the negative values of the signed distance function (SDF) of $\mathbf{P}^{\mathcal{T}}$ to $\mathbf{P}^{\mathcal{O}}$ to penalize any penetration between the robot hand and the object, and $\mathcal{L_{\ell}}$ computes the difference between two 6D poses:
@@ IV Experiments
  • Q7: Can our method generalize to novel robot hands? (See Appendix -B)
Figure 4: Visualization of all methods.
@@ IV-A Evaluation Metric
Success Rate: We evaluate the success of grasping by determining whether the force closure condition is satisfied. To implement this evaluation criterion, we used the Isaac Gym simulator [33]. A simple grasp controller is applied to execute the predicted grasps in the simulation (See Appendix -D). Certain forces are applied sequentially along six orthogonal directions, following the approach in [12]. We apply each force for a duration of 1 second. We consider the grasp successful if the object’s resultant displacement stays below 2 cm after applying the six directional forces.
@@ IV-C Overall Performance
Our experiments demonstrate that $\mathcal{D(R,O)}$ significantly outperformed all baselines regarding success rate across the robots by a large margin, highlighting the effectiveness of our approach. Fig. 4 visualizes grasps generated by our method alongside typical failure grasp poses from baselines. Our approach generates reasonable grasps, while DFC often results in unnatural poses. GenDexGrasp struggles with objects of complex shapes, frequently encountering significant penetration issues. Although ManiFM produces visually appealing grasps, its point-contact method lacks stability, lowering its success rate in simulation.

### v4

@@ I Introduction
Dexterous grasping is crucial in robotics as the first step in executing complex manipulation tasks. However, quickly obtaining a high-quality and diverse set of grasps remains challenging for dexterous robotic hands due to their high degrees of freedom and the complexities involved in achieving stable, precise grasps. Researchers have developed several optimization-based methods to address this challenge  [1, 2, 3, 4, 5]. Some of these methods, however, often focus on fingertip point contact, rely on complete object geometry, and require significant computational time to optimize. As a result, data-driven grasp generation methods have gained attention. These methods aim to solve the grasping problem using learning-based techniques. We can broadly categorize them into two types: those that utilize robot-centric representations, such as wrist poses and joint values [6, 7, 8], and those that rely on object-centric representations, such as contact points [9, 10, 11] or contact maps [12, 13, 14, 15].
Robot-centric representations (e.g., joint values), as used in methods like UniDexGrasp++ [8], directly map observation to control commands, enabling fast inference. However, these methods often exhibit low sample efficiency. Furthermore, these approaches struggle to generalize across different robotic embodiments, as the learned mappings are specific to the training data and do not quickly adapt to new robot designs or geometries. Object-centric representations (e.g., key points, contact points, affordances) effectively capture the geometry and contacts of objects, allowing for generalization across different shapes and robots, as demonstrated by methods like UniGrasp [9] and GenDexGrasp [12]. However, these methods are often less efficient as they typically require an additional optimization step—such as solving fingertip inverse kinematics (IK) or fitting the predicted contact maps under penetration-free and joint limit constraints to translate the object-centric representation into actionable robot commands. This optimization process is time-consuming due to its complexity and nonconvexity [16].
@@
Full-hand Contact
@@
Given the point clouds of both an open robotic hand and the object, our network predicts the $\mathcal{D(R,O)}$ representation, a matrix that encodes the relative distances between the point clouds of the object and the robotic hand in the desired grasping pose. Using this representation, we apply a multilateration method [17] to estimate the robot’s point cloud at the predicted pose, allowing us to compute the 6D pose of each hand link in the world frame and ultimately determine the joint configurations. To encode robotic hands, we propose a configuration-invariant pretraining method that learns the inherent alignment between various hand configurations, promoting grasp generation performance and cross-embodiment generalization. We validate the effectiveness of our approach through extensive experiments in both simulation and real-world settings. Our model achieves an average success rate of 87.53% in simulation across three dexterous robotic hands and an 89% success rate in real-robot experiments, demonstrating its robustness and versatility.
@@ II-A Learning-based Robotic Dexterous Grasping
Another line of research focuses on robot-centric approaches, which directly infer robot poses [18] or joint angles [8]. These methods are typically easier to execute as they directly infer joint values for dexterous hands. However, reinforcement learning in high-dimensional action spaces often suffers from sample inefficiency and is commonly trained in simulation. The sim-to-real gap [19, 20] further complicates policy transfer, with some works only reporting simulation results without real-world validation. To address these challenges, several methods leverage object-centric features [21], human grasp priors [22], or interaction bisector surface as the observation representation [23] to accelerate policy generation. Despite these efforts, the robot-centric approach limits the generalizability across different robotic embodiments. To combine the strengths of both approaches, we propose learning the relative distances as an interaction relationship between the robot and object for dexterous robotic grasping. This approach achieves robust real-world grasping performance and cross-embodiment generalization.
@@ III Method
III-A Configuration-Invariant Pretraining
Learning dexterous grasping involves understanding the spatial relationships between the robot hand and the object. The objective is to match the robot hand in a specific configuration with the object. However, this is challenging because the local geometric features of a point in the open-hand configuration may not align with those in the grasp configuration due to significant variations during articulation.
@@ III-A Configuration-Invariant Pretraining
Specifically, to establish correspondence between open-hand and close-hand configurations, we randomly sample a successful grasp $q_{\mathcal{A}}$ from the dataset and compute the corresponding canonical configuration $q_{\mathcal{B}}$ with a similar wrist pose. To align point clouds with the same point order, we uniformly sample points on the surface of each link for each robotic hand, storing the resulting point clouds as $\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}$, where $N_{\ell}$ is the number of links. We define a point cloud forward kinematics model, $\text{FK}\left(q,\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}\right)$ to map joint configurations to point clouds. Using this model, we obtain two point clouds $\mathbf{P}^{\mathcal{A}},\mathbf{P}^{\mathcal{B}}\in\mathbb{R}^{N_{\mathcal{R}}\times 3}$, representing these two joint configurations. Here, $N_{\mathcal{R}}$ is the number of points in the robot point cloud, set to 512 in practice.
@@ III-B $\mathcal{D(R,O)}$ Prediction
Given the wrist pose which can be either randomly sampled or user-specified, we obtain an open-hand configuration $q_{\text{init}}$. The robot point cloud under $q_{\text{init}}$ is $\mathbf{P}^{\mathcal{R}}=\text{FK}\left(q_{\text{init}},\left\{\mathbf{P}_{\ell_{i}}\right\}_{i=1}^{N_{\ell}}\right)\in\mathbb{R}^{N_{\mathcal{R}}\times 3}$, and the object point cloud is $\mathbf{P}^{\mathcal{O}}\in\mathbb{R}^{N_{\mathcal{O}}\times 3}$, where the number of points $N_{\mathcal{O}}$ is also 512 in practice. The objective of our neural network is to predict the point-to-point distance matrix $\mathcal{D(R,O)}\in\mathbb{R}^{N_{\mathcal{R}}\times N_{\mathcal{O}}}$, where both point clouds share the same origin.
Point Cloud Feature Extraction We begin by extracting point cloud embeddings using two encoders, $f_{\theta_{\mathcal{R}}}(\mathbf{P}^{\mathcal{R}})$ and $f_{\theta_{\mathcal{O}}}(\mathbf{P}^{\mathcal{O}})$, which share the same architecture. Specifically, we use a modified DGCNN [25] to better capture local structures and integrate global information (Appendix -F1). The robot encoder is initialized with pretrained parameters, using the method described in Sec. III-A, and remains frozen during training. These encoders extract point-wise features, $\bm{\phi}^{\mathcal{R}}$ and $\bm{\phi}^{\mathcal{O}}$ from the robot and object point clouds:
@@
To establish correspondences between the robot and object features, we apply two multi-head cross-attention transformers [26] (Appendix -F2), $g_{\theta_{\mathcal{R}}}(\bm{\phi}^{\mathcal{R}},\bm{\phi}^{\mathcal{O}})$ and $g_{\theta_{\mathcal{O}}}(\bm{\phi}^{\mathcal{O}},\bm{\phi}^{\mathcal{R}})$. These transformers integrate the relationships between the two feature sets, embedding correspondence information. This process maps the robot and object features to two sets of correlated features, $\bm{\psi}^{\mathcal{R}}$ and $\bm{\psi}^{\mathcal{O}}$:
@@
where $\sigma$ denotes the softplus function, and $\mathcal{N}_{\theta}$ is an MLP, which takes in the feature of $\mathbb{R}^{N_{\mathcal{O}}\times(2D+2d)}$ and outputs a positive number (Appendix -F3). By calculating on all $(\widehat{\bm{\psi}}_{i}^{\mathcal{R}},\widehat{\bm{\psi}}_{j}^{\mathcal{O}})$ pairs, we obtain the complete $\mathcal{D(R,O)}$ representation:
@@ III-C Grasp Configuration Generation from $\mathcal{D(R,O)}$
As shown in [29], this problem has a closed-form solution, and by using the implementation from [28], we can directly compute $p_{i}^{\prime\mathcal{R}}$. Repeating this process for each row of $\mathcal{D(R,O)}$ yields the complete predicted robot point cloud $\mathbf{P}^{\mathcal{P}}$ in the grasp pose. In 3D space, we can determine a point’s position by measuring its relative distances to just 4 other points. Our $\mathcal{D(R,O)}$ representation provides $N_{\mathcal{O}}(=512)$ relative distances, enhancing robustness to prediction errors.
@@
where $\mathbf{x}_{i}$ and $\mathbf{R}_{i}$ represent the translation and rotation of the $i$-th link, respectively. This computation can be directly performed using singular value decomposition (SVD).
Joint Configuration Optimization After predicting the 6D pose for each link, our objective is to optimize the joint values to align the translation of each link with the predicted result. During the inference phase, we initialize with $q_{init}$ and iteratively refine the solution through the following optimization problem using CVXPY [30]:
@@
In each iteration, the system computes the delta joint values $\delta\bm{q}$ by minimizing the objective function and updates the joint values as $\bm{q}\leftarrow\bm{q}+\delta\bm{q}$. Here, $\mathbf{x}_{i}$ represents the current link translation, $[\bm{q}_{min},\bm{q}_{max}]$ denotes the joint limits, and $\varepsilon_{q}=0.5$ is the maximum allowable step size. The optimization process can be efficiently parallelized, stably achieving convergence within one second, even for a 6+22 DoF ShadowHand.
@@ III-D Loss Function
where $\lambda_{\mathcal{D}}$, $\lambda_{\mathcal{T}}$, $\lambda_{\mathcal{P}}$, $\lambda_{KL}$ are hyperparameters for loss weights. The superscript “GT” refers to the ground truth annotations. $\mathcal{N}(0,I)$ is a standard Guassian distribution, and $\mathbf{P}^{\mathcal{T}}$ is the robot point cloud under the $\bm{\mathcal{T}^{*}}$ described in  III-C. $\mathcal{L_{\text{P}}}$ computes the sum of the negative values of the signed distance function (SDF) of $\mathbf{P}^{\mathcal{T}}$ to $\mathbf{P}^{\mathcal{O}}$ to penalize any penetration between the robot hand and the object, and $\mathcal{L_{\ell}}$ computes the difference between two 6D poses:
@@ IV Experiments
Figure 4: Visualization of generated grasps, compared to typical failure cases from existing approaches.
@@
  • Q7: Can our method generalize to novel robot hands? (Appendix -B)
@@ IV-A Evaluation Metric
Success Rate: We evaluate the success of grasping by determining whether the force closure condition is satisfied. To implement this evaluation criterion, we used the Isaac Gym simulator [31]. A simple grasp controller is applied to execute the predicted grasps in the simulation (Appendix -D). Following the metric in [12], we sequentially apply forces along six orthogonal directions, each for a duration of 1 second. The grasp is considered successful if the object’s resultant displacement remains below 2 cm after all six directional forces have been applied.
@@ IV-C Overall Performance
Our experiments demonstrate that $\mathcal{D(R,O)}$ significantly outperformed all baselines regarding success rate across the robots by a large margin, highlighting the effectiveness of our approach. For successful grasps of our method, the average displacement remains under 2 mm, with an average rotation below $1^{\circ}$, highlighting the firmness of our generated grasps. Fig. 4 visualizes grasps generated by our method alongside typical failure grasp poses from baselines. DFC often results in unnatural poses. GenDexGrasp struggles with objects of complex shapes, frequently encountering significant penetration issues. Although ManiFM produces visually appealing grasps, its point-contact method lacks stability, lowering its success rate in simulation.